Fermi National Laboratory
 

Accelerator Update for January 24, 2007 - January 26, 2007

 

Wednesday January 24

The day shift began with Operations monitoring the Tevatron (TeV) store 5194 (183.33E10) and the Recycler stash (50.9E10), with the Antiproton Source (Pbar) stacking (44.7E10), with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

A TeV expert began a Tevatron Electron Lens (TEL2) study at 10:10 AM.

A MI expert began a three-hour barrier bucket study at 11:33 AM.

At 12:41 PM, a Linac Klystron RF station (KRF3) tripped off. It reset okay.

Switchyard (SY) and Meson experts began commissioning MTest in the pion mode at 4 PM.

Operations reported at 4 PM that the Recycler had stashed antiprotons four times during the day shift.

At 4:14 PM, the NuMI beam permit tripped off. It reset okay.

At 4:44 PM, the MI beam permit tripped off for the first of four times due to bad beam positions. It reset each time.

The SY and Meson MTest pion mode commissioning ended at 5:41 PM.

A MI expert began a three-hour slipstacking study at 6:28 PM.

The TEL2 studies ended at 8:06 PM.

Operations reported at midnight that the Recycler had stashed four times during the day shift.

Thursday January 25

The midnight shift began with Operations monitoring store 5194 (57.01) and the stash (294.1E10), with Pbar just resuming to stack (0E10), and with NuMI taking beam.

At 12:22 AM, the Linac Klystron Debuncher tripped off on a cavity LCW flow indication. The cavity pump had tripped off. Everything reset.

At 12:27 AM, a Pbar Accumulator power supply (A:IB) tripped off on a ground fault. The stack was lost (41mA). Operators couldn’t reset the supply. Operations contacted an expert who said that operators should wait four hours and then call in a power supply expert.

Since Pbar wasn’t taking beam, Operations increased the beam to NuMI at 12:42 AM.

Operations prepared Pbar for access at 12:50 AM.

Operations turned off beam to Booster at 5 AM in preparation of today’s access.

Operations terminated store 5194 at 5:27 AM.

Operations prepared CDF and D0 for access at 5:49 AM.

Operations prepared NuMI for a MI-65 access at 6 AM.

Operations prepared Booster for access at 6:23 AM.

Operations reported at 8 AM that the Recycler stash was at 313.5E10. Experts and technicians had begun their accesses into the TeV, CDF and D0, NuMI, the Pbar rings and transport tunnel for maintenance work. An access into the MI-8 line would occur later this morning. This shutdown, which began around 6:30 AM, would last about ten hours.

Operations reported at 4 PM that the Recycler had held its stash. They reported a partial list of the day’s events: Interlocks were dropped in the Booster, CDF, and F-Sector. Techs rotated the Booster 95 LCW pumps at CUB. Operations rebooted many control computers. The Duty Electrician energized feeder 36 and partially restored power to Sites 50 & 52. HIL and PS-4 remained on generators.

Operations reported at 4:47 PM that they had begun to turn on the Proton Source and the TeV.

D0 completed their access at 5:04 PM.

At 5:05 PM, Recycler experts reported that a MI quadrupole power supply at MI-10 was having a problem at the end of flattop that in turn affected the Recycler quadrupole current compensation. The Recycler had lost about 30mA of its stash due to this problem. Operations called in a power supply expert.

At 5:39 PM, a Linac Klystron Debuncher trip off on a cavity LCW flow indication. This was the first of three trips over 22 minutes. It reset after each trip.

Operations established beam in the Booster at 5:54 PM.

CDF came out of access at 5:57 PM.

Operations reported at 6:59 PM that the MI ramp was on, the Recycler compensation was tracking, and the stash was stable. However, MECAR had caused a problem earlier that left a defocusing quadrupole in MI-10 in the wrong state. The MECAR problem was resolved.

At 7:32 PM, Operations reported that the Linac Klystron Debuncher was still having problems. Experts came in to investigate. This held off beam.

Operations put the TeV through a dry squeeze at 7:48 PM.

Linac experts completed their work on the Debuncher water skid. Operations established beam to NuMI at 9:45 PM.

A TeV expert reported at 11:29 PM that the new sextupole circuits were working well. The TeV was still in a dry squeeze to check this out.

Operations made the following Pbar progress report at midnight: The work in the transport line was completed at 8 PM. The nickel target replacement was finished by 9:30 PM. The magnet change was still underway at midnight.

Friday January 26

The midnight shift began with Operations monitoring the stash (279.8E10), with Pbar work continuing, and with NuMI taking beam.

At 12:08 AM, the NuMI beam permit tripped off due to beam position problems. It reset.

Operations put the TeV into shot setup around 2:30 AM.

Operators searched and secured the Pbar Rings at 3:09 AM.

At 4:02 AM, Operations reported that MECAR was still having problems, but this time it was affecting the horizontal quadrupole bus. Experts helped operators investigate the trouble. Operators removed and added events to the Time Line Generator (TLG).

Operations returned the Timeline module 130 to its original configuration (Load Collider Protons – Plus NuMI) and began loading protons into the TeV around 5:09 AM. Antiprotons were loaded at 5:19 AM.

At 5:59 AM, the TeV sector D17 horizontal separator sparked during the squeeze and took some luminosity away from the store.

Operations established store 5197 at 6:09 AM with an initial luminosity of 193.19E30.

The vacuum techs reported at 6:57 AM that the Pbar vacuum was still pumping down. It would be a few more hours before they could pull their portable vacuum pumps out of the tunnel.

The Plans for Friday and the Weekend

The plans for today are to finish the Pbar work and resume stacking, to troubleshoot MECAR, to establish beam to MTest and continue commission its beam, and to allow MI slipstacking and barrier bucket studies today and a barrier bucket study tomorrow. The plans for the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/29/2007    email Fermilab